A mass m = 2 kg is hung from a dynamometer whose upper end is fixed to the ceiling of an elevator. If the dynamometer reads a force P = 8 N. What can we say about the motion of the elevator?(Assume g = 10 m/s2)
A: the elevator has an acceleration a = 4 m/s2 directed downward
B: the elevator has an acceleration a = 6 m/s2 directed upward
C: the elevator has an acceleration a = 6 m/s2 directed downward
D: the elevator is moving upward
E: the elevator has an acceleration a = 4 m/s2 directed upward
1
Expert's answer
2018-06-04T08:58:08-0400
For this case:
ma=P-mg ,
where a- acceleration of the elevator (positive value means upward motion);
a=(P-mg)/m= (8-2∙10)/2= -6 m/〖s.〗^2

Answer: C: the elevator has an acceleration a = 6 m/s2 directed downward